You are given the following information about the demand for fresh tomatoes.

Using the midpoint method calculate the price elasticity of demand (using 2 decimal places) for the R30 to R32 price range. The price elasticity of demand for tomatoes is  

 


1
Expert's answer
2021-07-15T10:37:47-0400
"E=\frac{Q_2-Q_1}{Q_2+Q_1}\times \frac{p_2+p_1}{p_2-p_1}"